# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ III-B Upper Bound on Critical Threshold
[…]
    $\mathbb{P}_{H_{0}}\left(T\geq t+c_{6}\frac{n\|\pi\|_{\infty}^{2}}{k}\right)\leq\exp\bigg(-c_{6}^{\prime}\min\bigg{\{}\frac{2k(k-1)t^{2}}{nd_{\max}\|\pi\|_{\infty}^{4}},\frac{kt}{\|\pi\|_{\infty}^{2}}\bigg{\}}\bigg{missing})+\frac{\tilde{c}_{6}}{n^{3}}.$
@@
[…]
    $\mathbb{P}_{H_{1}}\left(T\leq\bigg{(}D-\frac{c_{7}n\|\pi\|_{\infty}}{\sqrt{nk}}\bigg{)}^{2}-t\right)\leq\exp\bigg(-c^{\prime}_{7}\min\bigg{\{}\frac{2k(k-1)t^{2}}{nd_{\max}\|\pi\|_{\infty}^{4}},\frac{kt}{\|\pi\|_{\infty}^{2}}\bigg{\}}\bigg{missing})+\exp\left(-c^{\prime\prime}_{7}\frac{kt^{2}}{\|\pi\|_{\infty}^{2}D^{2}}\right)+\frac{\tilde{c}_{7}}{n^{3}},$
@@
Theorem 3 established in Section V-D. It is worth mentioning that by Assumption 4, we have $\|\pi\|_{\infty}=\Theta(1/n)$, which causes the separation distance to scale like $O(1/\sqrt{nk})$ in order to achieve a minimax risk less than unity.
@@ III-D Lower Bound on Critical Threshold
In this section, we present an information-theoretic lower bound on the critical threshold for the hypothesis testing problem in Eq. 19 for the special case of a complete induced graph ${\mathcal{G}}$. Our lower bound demonstrates the minimax optimality of the scaling of the critical threshold provided in the upper bound in Theorem 2 for complete induced graphs. (We will show in Section III-E that pairwise comparison models on complete graphs can satisfy Assumptions 1, 2, 3 and 4.) As noted earlier, we assume that $k_{ij}=k$ for all $i,j\in[n]$ with $i\neq j$.
@@
The proof of Theorem 4 is provided in Section VI-A. The proof uses the Ingster-Suslina method for constructing a lower bound on the critical threshold [71]. This method is similar to the well-known Le Cam’s method, but it establishes a minimax lower bound by considering a point and a mixture on the parameter space instead of just two points. (Although Le Cam’s method could also be used for this proof in principle, the Ingster-Suslina method greatly simplifies the calculations to bound TV distance in our setting.) Our proof constructs a perturbed family of pairwise comparison models from a fixed BTL model and utilizes the complete graph structure to compute both the stationary distribution and the separation metric $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ in closed form under $H_{1}$. We remark that due to our problem setting, our proof here is much simpler than and quite different to the technique developed in [50], where the separation distance is quantified for Eulerian graphs in terms of sums of TV distances. We leave the problem of determining minimax lower bounds for more general graph topologies satisfying Assumptions 1, 2, 3 and 4 for future work. Finally, we also note that the bounds in Theorems 2 and 4 portray that the $\Theta(1/\sqrt{nk})$ scaling of the critical threshold is minimax optimal even if we took suprema over all induced graphs satisfying Assumptions 1, 2, 3 and 4 inside the infimum in the definition of minimax risk in Eq. 22.
@@ III-E Graphs with Bounded Principal Ratio
In this section, we establish bounds on both the principal ratio and edge expansion for three distinct classes of graphs: complete graphs, $\tilde{d}$-regular graphs, and random graphs generated from Erdős-Rényi models. These three classes represent a few examples of graphs for which Assumptions 2, 3 and 4 hold and the theoretical guarantees of our testing framework are valid for any pairwise comparison matrix $P$ consistent with Assumption 1. In the first case, we assume that the induced graph is complete and pairwise comparisons among all pairs are observed. The second scenario involves $\tilde{d}$-regular graphs that are sufficiently dense (as explained later) and possess some degree of edge expansion. And the third case assumes the existence of a complete underlying pairwise comparison model consistent with Assumption 1, where comparisons between any pair $(i,j)$ (and $(j,i)$), for $i>j$, are observed independently with probability $p$. We show that there exists a constant $\mathsf{c}_{p}>0$ such that as long as $np\geq\mathsf{c}_{p}\log n$, the Erdős-Rényi pairwise comparison model satisfies Assumptions 2, 3 and 4 with high probability. We also note that much of our technical analysis in each of the three scenarios lies in analyzing the principal ratio.
@@ III-E1 Complete Graphs
We begin by deriving bounds on the principal ratio of a canonical Markov matrix $S$ and the edge expansion of the DTM $R=\Pi^{1/2}S\Pi^{-1/2}$ corresponding to a complete graph. To this end, we consider a pairwise comparison matrix $P$ corresponding to a complete graph on $n$ vertices and construct $S$ using $d=2n$. (Note that $d=n$ would also work for the case of a complete graph.) In the following proposition, we show that the principal ratio is always upper bounded by $1/\delta^{2}$ for all pairwise comparison matrices consistent with Assumption 1.
Proposition 5 (Prinicipal Ratio for Complete Graphs).
@@
The proof is provided in Section A-A; our argument is a modification of that in [61, Theorem 3.1]. Proposition 5 illustrates that Assumption 4 holds with $h=1/\delta^{2}$. It also implies that the pairwise comparison model satisfies Assumption 2 with $\xi=\frac{\delta^{3}}{4(1+\delta)}$, because we obtain the following lower bound on the edge expansion $\phi(R)$ using Eq. 16 by substituting $d=2n$:
$\phi(R)\geq\frac{\delta^{3}}{2(1+\delta)n}\tilde{\phi}({\mathcal{G}})=\frac{\delta^{3}}{4(1+\delta)},$ (28)
where we have utilized the fact that for complete graphs, $\tilde{\phi}({\mathcal{G}})=\frac{n}{2}$. Alternatively, for complete graphs, we can obtain tighter upper bounds on the critical threshold (i.e., with better implicit dependence on $\delta$) by directly bounding the second largest singular value $\sigma_{2}(R)$ of the DTM $R$ (see Lemma 3 in Section V-A) instead of relying on expansion properties and Cheeger inequalities (i.e., using Assumption 2 and Lemma 4 in Section V-A). This alternative approach leverages bounds between contraction coefficients, specifically, in terms of the Dobrushin contraction coefficient for TV distance; see Section V-A for details. Additionally, note that a complete graph trivially satisfies Assumption 3. Thus, for the case of complete graphs, we have shown that any pairwise comparison model satisfying Assumption 1 also satisfies Assumptions 2, 3 and 4. This allows us to test whether data generated from any pairwise comparison matrix $P$ corresponding to a complete induced graph satisfying Assumption 1 conforms to an underlying BTL model. However, the testing procedure for a complete graph requires $n(n-1)k$ samples.
@@ III-E2 Dense $\tilde{d}$-Regular Graphs
Let $S$ be a canonical Markov matrix in Definition 2 corresponding to a $\tilde{d}$-regular graph ${\mathcal{G}}$ with $d=2\tilde{d}$ and stationary distribution $\pi$. Suppose further that Assumption 1 holds, and for some constants $a,b,c>0$, ${\mathcal{G}}$ satisfies $|{\mathcal{E}}(\mathcal{S},\mathcal{T})|\geq a|\mathcal{S}||\mathcal{T}|$ for all disjoint subsets $\mathcal{S},\mathcal{T}\subseteq[n]$ with $|\mathcal{S}|\geq bn$ and $|\mathcal{T}|\geq cn$. If $\tilde{d}\geq cn$ then, we have
@@
The proof is provided in Section A-B. This result generalizes the result in [72, Theorem 3], and illustrates that Assumption 4 holds with $h=\tilde{c}(a)/\delta^{5}$. The assumption, $|\mathcal{E}(\mathcal{S},\mathcal{T})|\geq a|\mathcal{S}||\mathcal{T}|$ for all disjoint subsets $\mathcal{S},\mathcal{T}$ with cardinality $\Theta(n)$, is typically satisfied by $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs ${\mathcal{G}}$ with $\tilde{d}=\ \nu n$ and $\lambda_{2}({\mathcal{G}})\leq(1-\tilde{\nu})\tilde{d}$ for some constants $\nu,\tilde{\nu}\in(0,1]$, where $\lambda_{2}({\mathcal{G}})$ denotes the second largest eigenvalue modulus of the adjacency matrix of ${\mathcal{G}}$; we refer readers to [73] for the definition of $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs. We also refer the readers to Section A-C for details regarding the existence of $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs with $\tilde{d}=\nu n$ and $\lambda_{2}({\mathcal{G}})\leq(1-\tilde{\nu})\tilde{d},$ and why they satisfy the assumption in Proposition 6. Furthermore, pairwise comparison models corresponding to such graphs also satisfy Assumption 2 as explained in Section A-C, cf. [74, Theorem 9.2.1], and trivially satisfy Assumption 3. Thus, for the case of dense $\tilde{d}$-regular graphs with appropriate expansion properties, we have again established that any pairwise comparison model satisfying Assumption 1 also satisfies Assumptions 2, 3 and 4. This permits us to test whether data generated from any pairwise comparison matrix $P$ corresponding to a dense $\tilde{d}$-regular graph satisfying Assumption 1 and certain expansion properties conforms to an underlying BTL model. Moreover, the denseness requirement $\tilde{d}=\Theta(n)$ implies that the testing procedure requires $\Theta(n^{2}k)$ samples in order to satisfy the various assumptions. The next setting establishes the utility of our testing procedure for sparse induced graphs.
@@ III-E3 Erdős-Rényi Random Graphs
In this case, we assume that there exists an underlying (pre-determined) pairwise comparison matrix $P^{*}$, corresponding to a complete graph, which satisfies Assumption 1. Subsequently, as in Erdős-Rényi random graphs [75], we sample each edge $(i,j)\in[n]^{2}$ with $i>j$ of the undirected induced graph independently with probability $p\in[0,1]$, such that for some sufficiently large constant $\mathsf{c}_{p}>1$, we have $np\geq\mathsf{c}_{p}\log n$ [^3] Note that the constant $\mathsf{c}_{p}$ does not depend on $p$. In essence, we randomly sample the entries of a fixed comparison matrix $P^{*}$ to obtain a pairwise comparison matrix $P={\mathcal{P}_{\mathcal{E}}}(P^{*})$, where ${\mathcal{E}}$ is the random edge set. Next, we construct canonical Markov matrices $S^{*}$ and $S$ corresponding to $P^{*}$ and $P$ with $d=3n$ and $d=3np$, respectively. The following well-known proposition provides high probability bounds on the vertex degrees for an Erdős-Rényi random graph, highlighting that $S$ is indeed a $(\frac{1}{2})$-lazy Markov matrix with high probability.
Proposition 7 (Degree Concentration for Erdős-Rényi Random Graph [11, Lemma 1]).
Suppose that the induced graph $\mathcal{G}$ is an Erdős-Rényi random graph with edges selected independently with probability $p$, and let $d_{\min}$ and $d_{\max}$ be the minimum and maximum degrees as defined in Assumption 3. If $p\geq\frac{c_{0}\log n}{n}$ for some sufficiently large constant $c_{0}>0$, then the event $\mathcal{A}_{0}$, defined as
@@
Let $\pi^{*}$ and $\pi$ be the stationary distributions of $S^{*}$ and $S$, respectively. Notably, by our assumption $np\geq\mathsf{c}_{p}\log n$ for $\mathsf{c}_{p}>1$, the underlying random graph ${\mathcal{G}}$ is connected with high probability [75]. Therefore, $S$ is irreducible and aperiodic with high probability (as $d_{\max}\leq 3np$ implies $S_{ii}>0$ for all $i\in[n]$), and hence, it has a unique stationary distribution $\pi$.
@@
Given an underlying pairwise comparison matrix $P^{*}$ with $n\geq 3$ satisfying Assumption 1, suppose we obtain the pairwise comparison matrix $P$ by randomly sampling the entries of $P^{*}$ according to an Erdős-Rényi model with parameter $p\in(0,1]$ satisfying $p\geq\frac{{c}_{9}\log n}{n}$ for some constant $c_{9}>1$. Then, there exists a constant $c_{10}>0$ such that with probability at least $1-O(n^{-3})$, the stationary distribution $\pi$ satisfies
@@
The proof of Theorem 5 is provided in Appendix B. The theorem immediately yields the following bound on the principal ratio corresponding to the sub-sampled pairwise comparison matrix $P$.
@@
Given an underlying pairwise comparison matrix $P^{*}$ with $n\geq 3$ satisfying Assumption 1, suppose we obtain the pairwise comparison matrix $P$ by randomly sampling the entries of $P^{*}$ according to an Erdős-Rényi model with parameter $p\in(0,1]$ satisfying $p\geq\frac{\mathsf{c}_{p}\log n}{n}$ for a constant $\mathsf{c}_{p}\geq\max\{c_{9},2c_{10}/\delta^{4}\}$ (where $c_{9},c_{10}$ are given in Theorem 5). Then, with probability at least $1-O(n^{-3})$, the principal ratio satisfies the bound
@@
The proof of Proposition 8 is provided in Section A-D. In particular, Proposition 8 illustrates that Assumption 4 holds with $h=7/\delta^{2}$ with high probability. Moreover, random graphs are known to have nice expansion properties [75, Theorem 2.8] (also see [76]). Indeed, using [77, Equation (11)], we have that for an Erdős-Rényi random graph, $\tilde{\phi}({\mathcal{G}})\geq\frac{1}{4}np$ with high probability, thereby showing that the pairwise comparison matrix $P$ also satisfies Assumption 2 with $\xi=\delta^{3}/(12\cdot 7\cdot(1+\delta))$ with high probability. Proposition 7 also implies that an Erdős-Rényi random graph satisfies Assumption 3 with high probability. Thus, for the case of Erdős-Rényi random graphs, we have again demonstrated that any pairwise comparison model satisfying Assumption 1 also satisfies Assumptions 2, 3 and 4 with high probability, as required for theoretical guarantees on our testing procedure. Notably, in this case, the testing procedure requires $O(kn\log n)$ pairwise comparisons, which matches the total number of observations needed for consistently estimating the parameters of the BTL model in [11].
III-F Stability of the BTL Assumption
In this section, we analyze the stability of the BTL assumption in the context of rankings for complete graphs. Specifically, we present an observation regarding how induced rankings behave when BTL models are perturbed by small amounts bounded by the scaling of the critical threshold of our testing framework. Recall that the BTL ranking orders agents based on the stationary distribution $\pi$ of the canonical Markov matrix (even for general pairwise comparison models, although usually, BTL rankings are used in the context of BTL models). Meanwhile, the Borda ranking is more general, as it does not rely on the BTL assumption and instead orders agents based on their Borda counts or scores [54] (defined next). We define the Borda count $\tau_{i}(P)$ of an agent $i\in[n]$ as the (scaled) probability that $i$ beats any other agent selected uniformly at random [54]:
$\tau_{i}(P)\triangleq\sum_{j=1}^{n}(1-p_{ij}).$ (29)
If the BTL assumption holds, then the Borda ranking equals the BTL ranking. Our goal here is to determine the size of the deviation from the BTL condition for a pairwise comparison model that is sufficient to produce a discrepancy between the BTL and Borda rankings. For simplicity in this section, we consider the symmetric setting where pairwise comparison matrices $P$ satisfy $p_{ij}+p_{ji}=1$ for all $i,j\in[n]$. The next proposition shows that the stability of the BTL assumption decreases as $n$ grows, i.e., as $n$ increases, smaller deviations from the BTL condition can lead to inconsistent BTL and Borda rankings.
Proposition 9 (Stability of BTL Assumption).
  1. 1.
    Define the error matrix $\mathsf{E}\triangleq\Pi P+P^{\mathrm{T}}\Pi-\mathbf{1}_{n}\pi^{\mathrm{T}}$. For any agents $i,j\in[n]$, $i\neq j$, the relative BTL and Borda rankings of $i$ and $j$ have the following relationship: $\pi_{i}\geq\pi_{j}$ if and only if
  2. 2.
    There exists a sub-sequence of pairwise comparison matrices $P\in(0,1)^{n\times n}$ such that agents $1$ and $2$ have constant Borda count difference $\Delta\tau\triangleq\tau_{1}(P)-\tau_{2}(P)>0$, but are ranked in the opposite order in a BTL ranking, i.e., $\pi_{1}<\pi_{2}$. [^4] We drop the subscript $n$ for the sub-sequence $P_{n}$ and write it as $P$ for clarity. Moreover, the deviation of $P$ from the BTL condition decays as
    for some constant $c_{11}>0$.
The proof is deferred to Section VI-B. Part 1 of Proposition 9 highlights the relationship between Borda and BTL rankings in terms of the weighted sum of entries of the error matrix. Moreover, part 2 conveys that the BTL assumption may potentially generate a “wrong” ranking (with respect to the Borda ranking) when the underlying pairwise comparison matrix is $O(1/\sqrt{n})$-separation distance away from the class of BTL models. Interestingly, this $O(1/\sqrt{n})$ deviation coincides with the critical threshold for the BTL testing problem (up to constant factors). It would be interesting to further explore the stability of the BTL assumption in the context of rankings in future work. We emphasize that our results here only consider stability in the complete graph case. When the induced graph is not complete, defining Borda rankings becomes ambiguous. Indeed, for a general graph, Borda rankings may not coincide with BTL rankings even under the BTL model.

### v2

@@ III-C Bounds on Type I and Type II Probabilities of Error
    There exist constants $c_{6},c^{\prime}_{6},\tilde{c}_{6}>0$ such that for all $t\geq 0$, and any BTL model in $\mathcal{M}_{0}$, the probability of type I error is upper bounded by
    $\mathbb{P}_{H_{0}}\bigg(T\geq u\frac{h^{2}}{nk}$ $+c_{6}\frac{h^{3}}{\xi^{4}}\frac{1}{nk}\bigg)\leq$
    $\exp\bigg(-c_{6}^{\prime}\min\bigg\{\frac{nu^{2}}{d_{\max}},nu\bigg\}\bigg.)+\frac{\tilde{c}_{6}}{n^{3}}.$
@@
    There exist constants $c_{7},c^{\prime}_{7},c^{\prime\prime}_{7},\tilde{c}_{7}>0$ such that for all $t\geq 0$, and any pairwise comparison model in $\mathcal{M}(\epsilon)$ with $\epsilon\geq c_{7}\sqrt{h^{3}}/\sqrt{nk\xi^{4}}$ the probability of type II error is upper bounded by
    $\mathbb{P}_{H_{1}}$ $\left(T\leq\bigg(D-\!{\frac{\sqrt{h^{3}}}{\xi^{2}}}\frac{c_{7}}{\sqrt{nk}}\bigg)^{2}\!\!-\frac{{u}}{nk}\right)\leq$
    $\ \ \ \ \ \ \ \ \ \ \ \ \ \exp\!\bigg(-c^{\prime}_{7}\min\bigg\{\frac{nu^{2}}{d_{\max}},\!nu\bigg\}\bigg)+$
    $\ \ \ \ \ \ \ \ \ \ \ \ \ \exp\left(-c^{\prime\prime}_{7}\frac{n^{2}\|\pi\|_{\infty}^{2}u^{2}}{D^{2}k}\right)+\frac{\tilde{c}_{7}}{n^{3}},$
@@
Theorem 3 is established in Section V-D.
@@ III-D Lower Bound on Critical Threshold
In this section, we present an information-theoretic lower bound on the critical threshold for the hypothesis testing problem in Eq. 17 for the special case of a complete induced graph $\mathcal{G}$}. Our lower bound demonstrates the minimax optimality of the scaling of the critical threshold provided in the upper bound in Theorem 2 for complete induced graphs. (We will show in Section III-E that pairwise comparison models on complete graphs can satisfy Assumptions 1 and 2.) As noted earlier, we assume that $k_{ij}=k$ for all $i,j\in[n]$ with $i\neq j$.
@@
The proof of Theorem 4 is provided in Section VI. The proof uses the Ingster-Suslina method for constructing a lower bound on the critical threshold [IngsterSuslina2003, RastogiBalakrishnanShahAarti2022]. This method is similar to the well-known Le Cam’s method, but it establishes a minimax lower bound by considering a point and a mixture on the parameter space instead of just two points. (Although Le Cam’s method could also be used for this proof in principle, the Ingster-Suslina method greatly simplifies the calculations to bound TV distance in our setting.) Our proof constructs a perturbed family of pairwise comparison models from a fixed BTL model and utilizes the complete graph structure to compute both the stationary distribution and the separation metric $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}({\mathbf{1}}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}/\|\pi\|_{\infty}$ in closed form under $H_{1}$. We remark that due to our problem setting, our proof here is much simpler than and different to the technique developed in [Seshadri2020], where the separation distance is quantified for Eulerian graphs in terms of sums of TV distances. We leave the problem of determining minimax lower bounds for more general graph topologies satisfying Assumptions 1 and 2 for future work. We also note that the bounds in Theorems 2 and 4 portray that the $\Theta(1/\sqrt{nk})$ scaling of the critical threshold is minimax optimal even if we took suprema over all induced graphs satisfying Assumptions 1 and 2 in the definition of minimax risk in Eq. 20. Finally, we remark that the construction of our lower bound inherently satisfies weak stochastic transitivity. Consequently, when the class of non-BTL models is restricted to the class of weakly stochastically transitive matrices, our lower bound result continues to hold. So, even under the restricted class of models satisfying weak stochastic transitivity, we obtain the same lower bound, and thus the testing problem is not simpler (in terms of critical threshold).
@@ III-E Graphs with Bounded Principal Ratio
In this section, we establish bounds on both the principal ratio and edge expansion for three distinct classes of graphs: complete graphs, $\tilde{d}-$regular graphs, and random graphs generated from Erdős-Rényi models. These three classes represent a few examples of graphs for which Assumptions 1 and 2 hold and the theoretical guarantees of our testing framework are valid for any pairwise comparison matrix $P$ consistent with Assumption 3. In the first case, we assume that the induced graph is complete and pairwise comparisons among all pairs are observed. The second scenario involves $\tilde{d}-$regular graphs that are sufficiently dense (as explained later) and possess some degree of edge expansion. And the third case assumes the existence of a complete underlying pairwise comparison model consistent with Assumption 3, where comparisons between any pair $(i,j)$ (and $(j,i)$), for $i>j$, are observed independently with probability pp. [^3] Throughout this work, $p$ without subscripts denotes the Erdős-Rényi edge probability, while $p_{ij}$ refers to pairwise comparison probabilities. We show that there exists a constant $\mathsf{c}_{p}>0$ such that as long as $np\geq\mathsf{c}_{p}\log n$, the Erdős-Rényi pairwise comparison model satisfies Assumptions 1 and 2 with high probability. We also note that much of our technical analysis in each of the three scenarios lies in analyzing the principal ratio.
@@ III-E1 Complete Graphs
We begin by deriving bounds on the principal ratio and the edge expansion of a canonical Markov matrix $S$ corresponding to a complete graph. To this end, we consider a pairwise comparison matrix $P$ corresponding to a complete graph on $n$ vertices and construct $S$ using $d=2n$. (Note that $d=n$ would also work for the case of a complete graph.) In the following proposition, we show that the principal ratio is always upper bounded by $1/\delta^{2}$ for all pairwise comparison matrices consistent with Assumption 3.
Proposition 5 (Principal Ratio for Complete Graphs).
@@
The proof is provided in Section A-A; our argument is a modification of that in [Minc1988, Theorem 3.1]. Proposition 5 illustrates that Assumption 2 holds with $h=1/\delta^{2}$. It also implies that the pairwise comparison model satisfies Assumption 1 with $\xi=\frac{\delta^{3}}{4(1+\delta)}$, because we obtain the following lower bound on the edge expansion $\phi(S)$ using Eq. 14 by substituting $d=2n$:
$\phi(S)\geq\frac{\delta^{3}}{2(1+\delta)n}\tilde{\phi}({\mathcal{G}})=\frac{\delta^{3}}{4(1+\delta)},$ (27)
where we have utilized the fact that for complete graphs, $\tilde{\phi}({\mathcal{G}})=\frac{n}{2}$. Alternatively, for complete graphs, we can obtain tighter upper bounds on the critical threshold (i.e., with better implicit dependence on $\delta$) by directly bounding the second largest singular value $\sigma_{2}(R)$ of the DTM $R=\Pi^{1/2}S\Pi^{-1/2}$ (see Lemma 3 in Section V-A) instead of relying on expansion properties and Cheeger inequalities (i.e., using Assumption 1 and Lemma 4 in Section V-A). This alternative approach leverages bounds between contraction coefficients, specifically, in terms of the Dobrushin contraction coefficient for TV distance (cf. [makur2019, MakurZheng2020]); see Section V-A for details. Thus, for the case of complete graphs, we have shown that any pairwise comparison model satisfying Assumption 3 also satisfies Assumptions 1 and 2. This allows us to test whether data generated from any pairwise comparison matrix $P$ corresponding to a complete induced graph satisfying Assumption 3 conforms to an underlying BTL model. However, the testing procedure for a complete graph requires n(n1)kn(n-1)k samples.
@@ III-E2 Dense $\tilde{d}-$Regular Graphs
Let $S$ be a canonical Markov matrix in Definition 2 corresponding to a $\tilde{d}-$regular graph $\mathcal{G}$} with $d=2\tilde{d}$ and stationary distribution $\pi$. Suppose further that Assumption 3 holds, and for some constants $a,b,c>0$, $\mathcal{G}$} satisfies $|{\mathcal{E}}(\mathcal{S},\mathcal{T})|\geq a|\mathcal{S}||\mathcal{T}|$ for all disjoint subsets $\mathcal{S},\mathcal{T}\subseteq[n]$ with $|\mathcal{S}|\geq bn$ and $|\mathcal{T}|\geq cn$. If $\tilde{d}\geq cn$, then we have
@@
The proof is provided in Section A-B. This result generalizes the result in [AksoyChung2016, Theorem 3], and illustrates that Assumption 2 holds with $h=\tilde{c}(a)/\delta^{5}$. The assumption, $|\mathcal{E}(\mathcal{S},\mathcal{T})|\geq a|\mathcal{S}||\mathcal{T}|$ for all disjoint subsets $\mathcal{S},\mathcal{T}$ with cardinality $\Theta(n)$, is typically satisfied by $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs $\mathcal{G}$} with $\tilde{d}=\ \nu n$ and $\lambda_{2}({\mathcal{G}})\leq(1-\tilde{\nu})\tilde{d}$ for some constants $\nu,\tilde{\nu}\in(0,1]$, where $\lambda_{2}({\mathcal{G}})$ denotes the second largest eigenvalue modulus of the adjacency matrix of $\mathcal{G}$}; we refer readers to [AlonChung1989] for the definition of $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs. We also refer the readers to Section A-C for details regarding the existence of $(n,\tilde{d},\lambda_{2}({\mathcal{G}}))$-regular expander graphs with $\tilde{d}=\nu n$ and $\lambda_{2}({\mathcal{G}})\leq(1-\tilde{\nu})\tilde{d}$, and why they satisfy the assumption in Proposition 6. Furthermore, pairwise comparison models corresponding to such graphs also satisfy Assumption 1 as explained in Section A-C, cf. [AlonSpencer2008, Theorem 9.2.1]. Thus, for the case of dense $\tilde{d}-$regular graphs with appropriate expansion properties, we have again established that any pairwise comparison model satisfying Assumption 3 also satisfies Assumptions 1 and 2. This permits us to test whether data generated from any pairwise comparison matrix $P$ corresponding to a dense $\tilde{d}-$regular graph satisfying Assumption 3 and certain expansion properties conforms to an underlying BTL model. Moreover, the denseness requirement $\tilde{d}=\Theta(n)$ implies that the testing procedure requires $\Theta(n^{2}k)$ samples in order to satisfy the various assumptions. The next setting establishes the utility of our testing procedure for sparse induced graphs.
@@ III-E3 Erdős-Rényi Random Graphs
In this case, we assume that there exists an underlying (predetermined) pairwise comparison matrix $P^{*}$, corresponding to a complete graph, which satisfies Assumption 3. Subsequently, as in Erdős-Rényi random graphs [BelaBollobas], we sample each edge $(i,j)\in[n]^{2}$ with $i>j$ of the undirected induced graph independently with probability $p\in[0,1]$, such that for some sufficiently large constant $\mathsf{c}_{p}>1$, we have $np\geq\mathsf{c}_{p}\log n$. [^4] Note that the constant $\mathsf{c}_{p}$ does not depend on $p$. In essence, we randomly sample the entries of a fixed comparison matrix $P^{*}$ to obtain a pairwise comparison matrix $P={\mathcal{P}_{\mathcal{E}}}(P^{*})$, where $\mathcal{E}$} is the random edge set. Next, we construct canonical Markov matrices $S^{*}$ and $S$ corresponding to $P^{*}$ and $P$ with $d=3n$ and $d=3np$, respectively. The following well-known proposition provides high probability bounds on the vertex degrees for an Erdős-Rényi random graph, highlighting that $S$ is indeed a $(\frac{1}{2})$-lazy Markov matrix with high probability.
Proposition 7 (Degree Concentration for Erdős-Rényi Random Graph [Chenetal2019, Lemma 1]).
[…]
@@
Let $\pi^{*}$ and $\pi$ be the stationary distributions of $S^{*}$ and $S$, respectively. Notably, by our assumption $np\geq\mathsf{c}_{p}\log n$ for $\mathsf{c}_{p}>1$, the underlying random graph $\mathcal{G}$} is connected with high probability [BelaBollobas]. Therefore, $S$ is irreducible and aperiodic with high probability (as $d_{\max}\leq 3np$ implies $S_{ii}>0$ for all $i\in[n]$), and hence, it has a unique stationary distribution $\pi$.
@@
Given an underlying pairwise comparison matrix $P^{*}$ satisfying Assumption 3 with $n\geq\tilde{c}_{9}$ for some constant $\tilde{c}_{9}>1$, suppose we obtain the pairwise comparison matrix $P$ by randomly sampling the entries of $P^{*}$ according to an Erdős-Rényi model with parameter $p\in(0,1]$ satisfying $p\geq\frac{{c}_{9}\log n}{n}$ for some constant $c_{9}>1$. Then, there exists a constant $c_{10}>0$ such that with probability at least $1-O(n^{-3})$, the stationary distribution $\pi$ satisfies
@@
The proof of Theorem 5 is provided in Appendix B and utilizes the leave-one-out technique in [Chenetal2019]. The theorem immediately yields the following bound on the principal ratio corresponding to the sub-sampled pairwise comparison matrix $P$.
@@
Given an underlying pairwise comparison matrix $P^{*}$ satisfying Assumption 3 with $n\geq\tilde{c}_{9}$, suppose we obtain the pairwise comparison matrix $P$ by randomly sampling the entries of $P^{*}$ according to an Erdős-Rényi model with parameter $p\in(0,1]$ satisfying $p\geq\frac{\mathsf{c}_{p}\log n}{n}$ for a constant $\mathsf{c}_{p}\geq\max\{c_{9},2c_{10}/\delta^{4}\}$ (where $c_{9},\tilde{c}_{9},c_{10}$ are given in Theorem 5). Then, with probability at least $1-O(n^{-3})$, the principal ratio satisfies the bound
@@
The proof of Proposition 8 is provided in Section A-D. In particular, Proposition 8 illustrates that Assumption 2 holds with $h=7/\delta^{2}$ with high probability. Moreover, random graphs are known to have nice expansion properties [BelaBollobas, Theorem 2.8] (also see [SipserSpielman1996]). Indeed, using [Han2023, Equation 11], we have that for an Erdős-Rényi random graph, $\tilde{\phi}({\mathcal{G}})\geq\frac{1}{4}np$ with high probability, thereby showing that the pairwise comparison matrix $P$ also satisfies Assumption 1 with $\xi=\delta^{3}/(12\cdot 7\cdot(1+\delta))$ with high probability. Thus, for the case of Erdős-Rényi random graphs, we have again demonstrated that any pairwise comparison model satisfying Assumption 3 also satisfies Assumptions 1 and 2 with high probability, as required for theoretical guarantees on our testing procedure. Notably, in this case, the testing procedure requires $O(kn\log n)$ pairwise comparisons, which matches the total number of observations needed for consistently estimating the parameters of the BTL model in [Chenetal2019].